# Maker-Breaker domination game played on corona products of graphs

### v2

@@ 2 Outcome of the game on corona products
In this section we determine the winner of the MBD game played on corona products. Before that we argue in the rest we can restrict our attention to corona products in which the first factor is of order at least two.
@@
Therefore, we may indeed restrict in the rest of the paper to corona products $G\odot H$, where $G$ is a graph with $n(G)\geq 2$. In what follows, we will also always assume that $V(G)=\{v_{1},\ldots,v_{n(G)}\}\subseteq V(G\odot H)$, and denote by $H_{i}$ the copy of $H$ in $G\odot H$ assigned to $v_{i}$, $i\in[n(G)]$.
@@
We now prove that the winner in the MBD game played on a corona product depends only on the winner of game played on the second factor.
@@ 3 MBD numbers of corona products
Proof. Using Theorem 1.3, a vertex $v_{i}$ of $G$ is an optimal first move for Dominator. Then all the vertices in its closed neighbourhood in $G\odot H$ are dominated. Staller responds by selecting a vertex $v_{j}$, where $j\neq i$. Staller’s plan afterwards is to play only on $H_{j}$. If Dominator plays his second move on $H_{j}$, then Staller can finish the game by selecting $\gamma_{\rm SMB}(H)$ vertices of $H_{j}$ and if Staller is able to play first in $H_{j}$, this can be done in $\gamma_{\rm SMB}^{\prime}(H)$ moves by Staller. As $\gamma_{\rm SMB}(H)\geq\gamma_{\rm SMB}^{\prime}(H)$, see [4, Proposition 2.7], Staller thus has a strategy to end the game in at most $1+\gamma_{\rm SMB}(H)$ moves. On the other hand, Dominator can force this number of moves by playing (optimally in the S-game on $H$) his second move in $H_{j}$, hence $\gamma_{\rm SMB}(G\odot H)\leq 1+\gamma_{\rm SMB}(H)$.
@@ 5 Concluding remarks
In this paper, we have explored the Maker-Breaker domination game on corona products $G\odot H$ of graphs. It is worth noting that this game is equivalent to the Maker-Breaker domination game played on the disjoint union of $n(G)$ copies of the graph obtained from $H$ by adding a new vertex adjacent to all the vertices of $H$. In other words, we are interested in the Maker-Breaker domination game played on the disjoint union of several copies of a given graph $H$ with $\gamma(H)=1$. Having this in mind it would be of interest to extend the results of this paper to generalized coronas which are defined just as coronas, except that the graphs associated to vertices of $G$ can be different. Generalized coronas were by now investigated from different angles, cf. [6, 19, 23].
For the the Maker-Breaker domination game on corona products, we have identified some general features, but several directions remain open for further exploration. Let us point to some of them. In Theorem 3.4 we have determined $\gamma_{\rm MB}(G\odot T)$ and $\gamma_{\rm MB}^{\prime}(G\odot T)$ for an arbitrary graph $G$ with with $n(G)\geq 2$ and an arbitrary tree $T$ with a perfect matching. It would be interesting to derive such a result also for arbitrary trees, that is, also for trees which do not contain perfect matchings. With respect to small MBD numbers it in particular remains to consider arbitrary graphs $G$ with $\gamma_{\rm SMB}(G)=3$, or with $\gamma_{\rm MB}(G)=3$, or with $\gamma_{\rm MB}^{\prime}(G)=3$, as well as corona products $G\odot H$ with $\gamma_{\rm MB}^{\prime}(G\odot H)=3$.

### v3

@@ 2 Outcome of the game on corona products
In this section we determine the winner of the MBD game played on corona products. Before that we argue that in the rest we can restrict our attention to corona products in which the first factor is of order at least two.
@@
Therefore, we may indeed restrict in the rest of the paper to corona products $G\odot H$, where $G$ is a graph with $n(G)\geq 2$, where $n(G)$ denotes the order of $G$. In what follows, we will also always assume that $V(G)=\{v_{1},\ldots,v_{n(G)}\}\subseteq V(G\odot H)$, and denote by $H_{i}$ the copy of $H$ in $G\odot H$ assigned to $v_{i}$, $i\in[n(G)]$.
@@
We now prove that the winner in the MBD game played on a corona product depends only on the winner of the game played on the second factor.
@@ 3 MBD numbers of corona products
Proof. Using Theorem 1.3, some vertex of $G$ is an optimal first move for Dominator, say $v_{i}$. Then all the vertices in its closed neighbourhood in $G\odot H$ are dominated. Staller responds by selecting a vertex $v_{j}$, where $j\neq i$. Staller’s plan afterwards is to play only on $H_{j}$. If Dominator plays his second move on $H_{j}$, then Staller can finish the game by selecting $\gamma_{\rm SMB}(H)$ vertices of $H_{j}$ and if Staller is able to play first in $H_{j}$, this can be done in $\gamma_{\rm SMB}^{\prime}(H)$ moves by Staller. As $\gamma_{\rm SMB}(H)\geq\gamma_{\rm SMB}^{\prime}(H)$, see [4, Proposition 2.7], Staller thus has a strategy to end the game in at most $1+\gamma_{\rm SMB}(H)$ moves. On the other hand, Dominator can force this number of moves by playing (optimally in the S-game on $H$) his second move in $H_{j}$, hence $\gamma_{\rm SMB}(G\odot H)\leq 1+\gamma_{\rm SMB}(H)$.
@@ 5 Concluding remarks
In this paper, we have explored the Maker-Breaker domination game on corona products $G\odot H$ of graphs. It is worth noting the following equivalent approach to it. Let $\widehat{H}$ be the graph obtained from $H$ by adding a new vertex adjacent to all the vertices of $H$, and let $k\widehat{H}$ denote the disjoint union of $k$ copies of $\widehat{H}$. Then the Maker-Breaker domination game played on the corona product $G\odot H$ is equivalent to the Maker-Breaker domination game played on $n(G)\widehat{H}$.
Clearly, $\gamma(\widehat{H})=1$. Having in mind the above equivalent formulation of the Maker-Breaker domination game on corona products, it would be of interest to extend the results of this paper to generalized coronas which are defined just as coronas, except that the graphs associated to vertices of $G$ can be different. Generalized coronas were by now investigated from different angles, cf. [6, 19, 23].
For the the Maker-Breaker domination game on corona products, we have identified some general features, but several directions remain open for further exploration. Let us point to some of them.
In Theorem 3.2 we have bounded the values of $\gamma_{\rm MB}(G\odot H)$ and of $\gamma_{\rm MB}^{\prime}(G\odot H)$ for the case when $o(H)=\mathcal{D}$. It would be interesting to investigate possible values that $\gamma_{\rm MB}^{\prime}(G\odot H)-\gamma_{\rm MB}(G\odot H)$ can attain.
In Theorem 3.4 we have determined $\gamma_{\rm MB}(G\odot T)$ and $\gamma_{\rm MB}^{\prime}(G\odot T)$ for an arbitrary graph $G$ with $n(G)\geq 2$ and an arbitrary tree $T$ with a perfect matching. It would be interesting to derive such a result also for arbitrary trees, that is, also for trees which do not contain perfect matchings.
With respect to small MBD numbers it in particular remains to consider arbitrary graphs $G$ with $\gamma_{\rm SMB}(G)=3$, or with $\gamma_{\rm MB}(G)=3$, or with $\gamma_{\rm MB}^{\prime}(G)=3$, as well as corona products $G\odot H$ with $\gamma_{\rm MB}^{\prime}(G\odot H)=3$.